A rectangle has a diagonal of 15.4cm and the width is 6.4cm. To the nearest centimeter what is the length of the rectangle?
Answer by JulietG(1812)

A rectangle split down the diagonal gives us two right triangles, so we can use the right triangle formula.
a^2 + b^2 = c^2
We're given c^2 as 15.4cm
We're given b^2 as 6.4cm (or we could call it a^2; works either way)
.
a^2 + 6.4^2 = 15.4^2
a^2 + 40.96 = 237.16
Subtract 40.96 from each side
a^2 = 196.2
Take the square root of each side
a = 14.0071
to the nearest cm
a = 14 <<--